OpenStudy (anonymous):

a 400 g bead slides on a curved frictionless wire, starting from rest at point a. find the speed of the bead at point b and point c.

OpenStudy (anonymous):



OpenStudy (anonymous):

attempt to the solution; m = 400 g = 0.4 kg A to B (1/2)(0.4)(0) +(0.4)(9.8)(5)=(1/2)(0.4)(V_B^2)+(0.4)(9.8)(0) V_B = 9.9 m/s ?? B to C (1/2)(0.4)(9.9^2)+0=(1/2)(0.4)(V_C^2)+(0.4)(9.8)(2) V_C = 7.67 m/s ??

OpenStudy (anonymous):

@zzr0ck3r @heedcom

OpenStudy (anonymous):

@AbhimanyuPudi

OpenStudy (anonymous):

I did not check the calculation part but the approach is absolutely right...conservation of energy

OpenStudy (anonymous):

thanks again! ..:)

OpenStudy (anonymous):

welcome : )